OpenStudy (anonymous):

At most how many acute angles can a convex 2012-gon have?

OpenStudy (cwrw238):

good - question - i would say none

jimthompson5910 (jim_thompson5910):

If you're talking about interior angles, then yes, the answer is none. This is because the measure of each interior angle x is given by the formula x = (180*(n-2))/n where n is the number of sides. If you plug in n = 2012, then x = (180*(2012-2))/2012 x = 179.821 approximately which is larger than 90 degrees. So all the interior angles are obtuse.

OpenStudy (cwrw238):

if it was a regular polygon each external angle would be 360 / 2112 degrees so internal angle is > 179 degrees

OpenStudy (anonymous):

ans is 3